OpenStudy (anonymous):

I need help really really fast for this radical expression pretty simple, will fan and medal!

OpenStudy (anonymous):

\[\sqrt{250h^4k^5}\]

OpenStudy (anonymous):

what is the expression?

OpenStudy (anonymous):

wht help

OpenStudy (anonymous):

right up there

OpenStudy (anonymous):

? what is that

OpenStudy (anonymous):

you just took out the 250 that doesnt really make sense

OpenStudy (anonymous):

5h^2k^2√10k this is right

OpenStudy (anonymous):

Thanks a lot taht was one of the answer choices so i beleive you

OpenStudy (anonymous):

250 * h^4 * k^5. If you take the square root of each of those terms and multiply you will get your answer. \[\sqrt{250} = 5 * \sqrt{10}\] \[\sqrt{h^4} = h^2\] \[\sqrt{k^5} = k^2\sqrt{k}\]

OpenStudy (anonymous):

So if you multiply that together it is \[h^2 * k^2 * 5 \sqrt{10k}\]